1.list the angles of triangle KLM with vertices K (-2, -2) , L (2,6), M (7, -2) in order from smallest to largest?

2.determine the perimeter of a square with diagonal of square root 72 centimeters.

Answer by checkley77(12569)

1). K (-2, -2) , L (2,6), M (7, -2)
Plotting these 3 points we get a tringle with the following angles:
K=64 degrees. The largest.
L=48 degrees. The smallest.
M=62 degrees. The middle.
2). Let x be the length of the sides of the square.
The diagonal is the hypotenuse of a right triangle with sides of x, x, & an hypotenuse of sqrt72.
x^2+x^2=sqrt72
2x^2=8.48
x^2=8.48/2
x^2=4.24
x=sqrt4.24
x=2.06 answer for the length of the sides of the square.
proof:
2.06^2+2.06^2=sqrt72
4.24+4.24=8.48
8.48=8.48